OpenStudy (anonymous):

What is the value of the expression 1/4^-3 ? A. 12 B. 64 C. 1/64 D. 1/12

OpenStudy (anonymous):

@domebotnos

OpenStudy (anonymous):

is it 64?

jimthompson5910 (jim_thompson5910):

Similar Problem \[\Large \frac{1}{2^{-4}} = \frac{2^{4}}{1}\] \[\Large \frac{1}{2^{-4}} = \frac{16}{1}\] \[\Large \frac{1}{2^{-4}} = 16\]

jimthompson5910 (jim_thompson5910):

I'm using the rule \[\Large \frac{1}{x^{-k}} = \frac{x^{k}}{1}\]

OpenStudy (anonymous):

is it 64?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thank You!

jimthompson5910 (jim_thompson5910):

np and nice work

OpenStudy (anonymous):

What is the value of the expression of 3/3^-2 i keep getting 1 for some reason

jimthompson5910 (jim_thompson5910):

that's because you evaluated 3/3

jimthompson5910 (jim_thompson5910):

but PEMDAS says that 3^(-2) must be evaluated first

jimthompson5910 (jim_thompson5910):

You can use that rule I posted to get \[\Large \frac{3}{3^{-2}} = 3*3^{2} = ???\]

OpenStudy (anonymous):

27?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thank You!